# supplement

### v1

@@ Abstract
This paper presents several forecasting methods to model and forecast subnational age distribution of death counts. The age distribution of death counts has many similarities to probability density functions, which are nonnegative and have a constrained integral, and thus live in a constrained nonlinear space. To address the nonlinear nature of objects, we implement a cumulative distribution function transformation that has an additional monotonicity. Using the Japanese subnational life-table death counts obtained from the Japanese Mortality Database (2025), we evaluate the forecast accuracy of the transformation and forecasting methods. The improved forecast accuracy of life-table death counts implemented here will be of great interest to demographers in estimating regional age-specific survival probabilities and life expectancy.
Keywords: constrained functional time series; cumulative distribution function transformation; (split) conformal prediction; life-table death counts; principal component analysis
@@ 1 Introduction
Density-valued objects for multiple groups are common, with examples including income distributions across different populations (Kneip & Utikal, 2001), financial return distributions for multiple stocks (Petersen et al., 2022), distributions of bidding times in online auctions for various items (Wang et al., 2008), and subnational age distributions of deaths in demography (Jiménez-Varón et al., 2025), among others. We present several forecasting methods to model and predict the subnational age distribution of deaths, which are also useful for applications in actuarial science, such as fixed-term or lifetime annuity pricing.
@@
Since age distributions of deaths naturally take the form of density functions, they are well-suited for modeling using advanced forecasting techniques from compositional and functional data analyses. Oeppen (2008) demonstrates that using compositional data analysis to forecast the age distribution of death counts performs similarly in accuracy to forecasting age-specific mortality rates. Within the compositional data analysis, the Lee-Carter model is frequently applied to model and forecast unconstrained data. Shang & Haberman (2020) extend this approach by incorporating a functional data analytic perspective, which includes multiple principal components and nonparametric smoothing. Recently, Shang & Haberman (2025) introduce a cumulative distribution function (CDF) transformation, converting each year’s age distribution of deaths into a probability bounded within the unit interval to address the presence of zero values. Through cumulative summation, a probability density function can be transformed into a CDF, which provides the additional benefit of monotonicity (see also Mayhew & Smith, 2013). With a time series of CDFs, it is common to model and extrapolate its pattern via a logistic transformation.
@@
Multiple density-valued objects are observed over time are related to high-dimensional functional time series (HDFTS). In the statistical literature, Zhou & Dette (2023) derive Gaussian and multiplier bootstrap approximations for sums of HDFTS. With these approximations, they construct joint simultaneous confidence bands for the mean functions and develop a hypothesis to test whether the mean functions in the cross-sectional dimension exhibit parallel behavior. Hallin et al. (2023) investigate the representation of HDFTS using a factor model, determining conditions on the eigenvalues of the covariance operator crucial for establishing the existence and uniqueness of the factor model. Gao et al. (2019) adopt a two-stage approach, combining truncated principal component analysis and a separate scalar factor model for the resulting panels of scores. Tavakoli et al. (2023) introduce a functional factor model with a functional factor loading and a vector of real-valued factors, while Guo et al. (2024) propose a functional factor model with a real-valued factor loading and a functional factor. Leng et al. (2024) introduce a unified framework that accommodates both types of factor models. Tang et al. (2022) study clustering for age-specific subnational mortality rates, which is an example of HDFTS. Li et al. (2024) introduce hypothesis tests and estimation procedures for testing and locating (common) change points. Chang et al. (2025) develop a two-stage procedure for modeling and forecasting HDFTS.
@@
Our contributions are twofold. First, we present several visualization techniques to capture the patterns in age distribution of death counts and display contrasts between subnational and national data. Second, we revisit several forecasting methods and compare their point and interval forecast accuracy with applications to subnational life-table death counts.
@@
Our paper is structured as follows. In Section 2, we describe our motivating data set and introduce a series of image plots for displaying important features in HDFTS. In Section 3, we present the CDF transformations. Within the CDF transformation, we consider a number of forecasting methods in Section 4 to model and forecast unconstrained functional time-series data. In Section 5, we propose two general strategies for constructing pointwise prediction intervals for the age distribution of death counts. In Section 6, we evaluate and compare point forecast accuracy using the Kullback-Leibler divergence (KLD) and Jensen-Shannon divergence (JSD), and interval forecast accuracy using the coverage probability difference (CPD) and interval score. Section 7 concludes with some ideas on how the methodology presented can be further extended.
@@ 2 Subnational age distribution of death in Japan
In many developed countries such as Japan, increases in longevity and an aging population have led to concerns about the sustainability of pensions, health, and aged care systems (see, e.g., Coulmas, 2007). These concerns have resulted in a surge of interest among government policy makers in accurately modeling and forecasting age-specific mortality. Subnational forecasts of age-specific mortality are useful for informing policy within local regions and any improvement in forecast accuracy of mortality is beneficial for determining the allocation of current and future resources at the national and subnational levels.
In demography, the age distribution of deaths provides important insights into longevity and lifespan variability that cannot be grasped directly from the central mortality rate or survival function. As pointed out by Oeppen (2008), the age distribution of death counts is more suitable than central mortality rates for computing life expectancy and annuity premia. In addition to providing an informative description of the mortality experience of a population, life-table death counts yield readily available information on the “central longevity indicators” (e.g., mean, median, and modal age at death, see Cheung et al., 2005; Canudas-Romo, 2010) and lifespan variability (e.g., van Raalte & Caswell, 2013; Aburto & van Raalte, 2018).
We study Japanese lifetable death counts from 1975 to 2022, obtained from the Japanese Mortality Database (2025). We consider ages from 0 to 110 in single years of age. Figure 1 shows rainbow plots of the female and male age-specific life-table death counts in Japan from 1975 to 2022. The time ordering of the curves follows the color order of a rainbow, where curves from the distant past are shown in red and the more recent curves are shown in violet. The figures show typical mortality curves for a developed country, with a decreasing trend in infant death counts. A typical negatively skewed distribution for the life-table death counts is apparent where the peaks shift to higher ages for both females and males. This shift is a main source of longevity risk, which is a major issue for insurers and pension funds, especially in the selling and risk management of annuity products (see Denuit et al., 2007, for a discussion). Moreover, the spread of the distribution indicates lifespan variability. A decrease in variability over time can be observed directly and can be measured by the Gini coefficient (Wilmoth & Horiuchi, 1999; Debón et al., 2017).
@@
Figure 1: Functional time series graphical displays of age-specific life-table death counts from 1975 to 2022 in a single-year group. The oldest years are shown in red, with the most recent years in violet. Curves are ordered chronologically according to the colors of the rainbow.
@@ 2.1 Image plots
Another visual perspective of the data is shown in the image plot of Figure 2. Here, we graph the symmetric KLD between life-table death counts for each prefecture to life-table death counts for the whole country, thus allowing relative mortality comparisons to be made. A divergent color palette is used with red representing larger KLD values (i.e., greater difference between subnational and national data) and yellow denoting smaller KLD values. The prefectures are ordered geographically from north to south, so the most northerly prefecture (Hokkaido) is at the top of the panels, and the most southerly prefecture (Okinawa) is at the bottom.
Figure 2: Image plots showing the symmetric KLD introduced in (9) between life-table death counts of a prefecture and life-table death counts of the whole country. The top panel shows mortality averaged over ages, while the bottom panel shows mortality averaged over years. Prefectures are numbered geographically from north to south.
The top row of the panels shows mortality for each prefecture and year, averaged over all ages. In 2011, in Iwate, Miyagi, and Fukushima, there was a large increase in mortality compared to other prefectures. These are northern coastal regions, and the inflated relative mortality is due to the tsunami of 11 March 2011 (see, e.g., Nakahara & Ichikawa, 2013) caused by a magnitude–9.0 earthquake.
@@
The bottom of the panels shows mortality for each prefecture and age, averaged over all years. Several striking features are apparent. There are strong differences between the prefectures for children and teenagers; this is possibly due to socio-economic differences and the accessibility of health services.
@@ 2.2 Functional cross-correlation plot
Let $d_{t,s}^{g}(u)$ be the age distribution of deaths for age $u$, prefecture $s$, gender $g$ at year $t$, respectively. There are 111 ages, $S=47$ prefectures, and $48$ years. Following Kokoszka et al. (2017) and Mestre et al. (2021), the functional autocorrelation coefficient at lag $h$ can be defined as
[…]
@@
[…]
@@
[…]
@@
In Figure 3, we display the functional cross-correlation functions showing the dependence of life-table death counts between each prefecture and the whole country. Each subnational data has a comparably stronger cross correlation with the national data at lags 1 and 2. The cross-correlation reduces as lag increases.
@@ 3 Cumulative distribution function transformation
In the CDF transformation, we first convert an age distribution of death for each year to a probability by dividing its radix $10^{5}$. Via the cumulative sum, we transform a probability density function into a CDF,
@@
where $D_{t,s}^{g}(u_{111})=1$. In doing so, it enjoys an additional benefit of monotonicity (see also Mayhew & Smith, 2013).
@@
where $\Delta$ represents the first-order differencing, where $\Delta_{i=1}^{1}D_{t,s}^{g}(u_{1})=D_{t,s}^{g}(u_{1})$. The constant $10^{5}$ is the radix of the life-table death count. Since the data are observed discretely, we use the cumulative sum and first-order differencing as approximations for integration and differentiation, respectively.
@@ 4 Functional time-series forecasting models
[…]

### v2

# Modeling and forecasting subnational age distribution of death counts

@@ Abstract
Existing mortality forecasting methods focus on age-specific mortality rates, which lie in an unconstrained space and overlook the distributional nature of life-table death counts. Few studies have developed and compared forecasting methods that model the shape and dynamics of the age distribution of deaths, especially at the subnational level, where data quality varies greatly. This paper presents several forecasting methods to model and forecast the subnational age distribution of death counts. The age distribution of death counts has many similarities to probability density functions, which are nonnegative and have a constrained integral, and thus live in a constrained nonlinear space. To address the nonlinear nature of objects, we implement a cumulative distribution function transformation that is scale-free and has additional monotonicity. Using subnational Japanese life-table death counts from Japanese Mortality Database (2025), we evaluate the forecast accuracy of the transformation and forecasting methods. The improved forecast accuracy of life-table death counts implemented here will be of great interest to demographers in estimating regional age-specific survival probabilities and life expectancy, and to actuaries for determining annuity prices for various ages and maturities.
Keywords: conformal prediction intervals; constrained functional time series; cumulative distribution function transformation; life-table death counts; principal component analysis
@@ 1 Introduction
Accurate prediction of subnational mortality rates plays a crucial role in both actuarial studies and demographic planning. From a demographic perspective, subnational mortality forecasts inform regional health planning, pension sustainability assessments, and population projections that guide policy decisions on healthcare infrastructure and resource allocation (see, e.g., Booth & Tickle, 2008). In actuarial science, understanding spatial variations in mortality and their evolution over time is essential for pricing life insurance products, designing annuities, and managing longevity risk in pension schemes (e.g., Cairns et al., 2006; Haberman & Renshaw, 2011). Subnational mortality prediction allows actuaries and policymakers to capture heterogeneity across regions, reflecting socioeconomic, environmental, and healthcare disparities, improving the reliability of longevity projections (Li & Lee, 2005).
Although there exist many multipopulation techniques, such as the Li & Lee’s (2005) method, for modeling and forecasting age-specific mortality rates, the age distribution of death counts is another mortality instrument, which possesses its unique advantages over mortality rates. For example, the age distribution of death counts preserves distributional information: the age distribution of deaths retains the shape of how deaths are spread across ages. Due to non-negativity and summability, the changes are relative in age. It captures shifts in longevity more directly: the mode and spread of the distribution can be measured. It is less sensitive to population size estimates; at the subnational level, these estimates may be unreliable. Although mortality rates at older ages can be outside their plausible range (0,1), it does not occur for the age distribution of death counts. It facilitates comparison across populations and time, as it is insensitive to population size estimates. The age distribution of death counts is an explicit way of measuring the probability of dying, commonly studied in the Cairns et al.’s (2006) model.
@@
Following the early works of Lu et al. (2020 and 2021) on subnational mortality rates and Bergeron-Boucher et al. (2018) on subnational age distribution of death counts, we present several forecasting methods to model and predict the subnational age distribution of deaths, which are also useful for applications in actuarial science, such as fixed-term or lifetime annuity pricing (see, e.g., Shang & Haberman, 2020). Apart from subnational age distributions of deaths in demography (Jiménez-Varón et al., 2025), density-valued objects for multiple groups are common, with examples including income distributions across different populations (Kneip & Utikal, 2001), financial return distributions for multiple stocks (Petersen et al., 2022), distributions of bidding times in online auctions for various items (Wang et al., 2008), among others.
@@
Since age distributions of deaths naturally take the form of density functions, they are well-suited for modeling using advanced forecasting techniques from compositional and functional data analyses. Oeppen (2008) demonstrates that the use of compositional data analysis to forecast the age distribution of death counts performs similarly in accuracy to forecast age-specific mortality rates. Within compositional data analysis, the Lee-Carter model is frequently applied to model and forecast unconstrained data. Shang & Haberman (2020) extend this approach by incorporating a functional data-analytic perspective, including multiple principal components and nonparametric smoothing. Recently, Shang & Haberman (2025b) introduce a cumulative distribution function (CDF) transformation, converting each year’s age distribution of deaths into a probability bounded within the unit interval to address the presence of zero values. Through cumulative summation, a probability density function can be transformed into a CDF, providing the additional benefit of monotonicity (see also Mayhew & Smith, 2013). With a time series of CDFs, it is common to model and extrapolate the pattern using a logistic transformation.
@@
Multiple density-valued objects observed over time are related to high-dimensional functional time series (HDFTS). In the statistical literature, Zhou & Dette (2023) derive Gaussian and multiplier bootstrap approximations for the sums of HDFTS. With these approximations, they construct joint simultaneous confidence bands for the mean functions and develop a hypothesis to test whether the mean functions in the cross-sectional dimension exhibit parallel behavior. Hallin et al. (2023) investigate the representation of HDFTS using a factor model and determine conditions on the eigenvalues of the covariance operator that are crucial for establishing the existence and uniqueness of the factor model. Gao et al. (2019) adopt a two-stage approach, combining truncated principal component analysis and a separate scalar factor model for the resulting panels of scores. Tavakoli et al. (2023) introduce a functional factor model with a functional factor loading and a vector of real-valued factors, while Guo et al. (2025) propose a functional factor model with a real-valued factor loading and a functional factor. Leng et al. (2026) introduce a unified framework that accommodates both types of factor models. Tang et al. (2022) study clustering for age-specific subnational mortality rates, which is an example of HDFTS. Li et al. (2024) introduce hypothesis tests and estimation procedures for testing and locating (common) change points. Chang et al. (2025) develop a two-stage procedure for modeling and forecasting HDFTS.
@@
Although numerous studies have addressed the modelling and forecasting of age distributions of death counts at the national level (see, e.g., Shang, 2025; Shang & Haberman, 2025a and b), there remains a notable gap in extending such forecasts to the subnational level. Our contributions are twofold. First, we present several visualization techniques to capture patterns in the age distribution of death counts and to display contrasts between subnational and national data. Second, we revisit several forecasting methods and compare their point and interval forecast accuracy with applications to subnational life-table death counts. For different sexes and prefectures, we provide some recommendations on the optimal multipopulation models, which can vary in terms of their point and interval forecast accuracies. The improvement in forecast accuracy of life-table death counts may translate to more accurate annuity prices and life expectancy calculations; in turn, these inform policymakers on the sustainability of the pension system and aged care facilities.
@@
Our paper is structured as follows. In Section 2, we describe our motivating data set and introduce a series of image plots for displaying important features in HDFTS. In Section 3, we present the CDF transformations. Within the CDF transformation, we consider a number of forecasting methods in Section 4 to model and forecast unconstrained functional time-series data. In Section 5, we propose two general strategies for constructing pointwise prediction intervals for the age distribution of death counts. In Section 6, we evaluate and compare point forecast accuracy using the Kullback-Leibler divergence (KLD) and Jensen-Shannon divergence (JSD), and interval forecast accuracy using the coverage probability difference (CPD) and mean interval score. In Section 6.6, we show relationship between life-table death count and survival probability and life expectancy. In Section 6.7, we present a model diagnostic analysis using functional analysis of variance. Section 7 concludes with ideas on how the methodology presented can be extended.
@@ 2 Subnational age distribution of death in Japan
In many developed countries, such as Japan, increases in longevity and an aging population have led to concerns about the sustainability of pension, health, and aged care systems (see, e.g., Coulmas, 2007). These concerns have led to a surge of interest among government policymakers in accurately modeling and forecasting age-specific mortality. Subnational forecasts of age-specific mortality are useful for informing policy at the local level, and any improvement in forecast accuracy is beneficial for allocating current and future resources at national and subnational levels.
In demography, the age distribution of deaths provides important insights into longevity and lifespan variability that cannot be grasped directly from the central mortality rate or survival function. As pointed out by Oeppen (2008), the age distribution of death counts is more suitable than central mortality rates for computing life expectancy and annuity premia. In addition to providing an informative description of the mortality experience of a population, life-table death counts yield readily available information on the “central longevity indicators” (e.g., mean, median, and modal age at death, see Cheung et al., 2005; Canudas-Romo, 2010) and lifespan variability (e.g., van Raalte & Caswell, 2013; Aburto & van Raalte, 2018). The life-table death count is less sensitive to population size estimate, especially at the subnational level; population size estimate may be unreliable. Because of this, it is insensitive to population size estimates, which facilitates comparisons across populations and over time. The life-table death count is directly linked to the probability of dying $q_{x}$, which is the mortality instrument in the Cairns et al.’s (2006) model.
We study Japanese life-table death counts from 1975 to 2023, obtained from the Japanese Mortality Database (2025). We observe complete life-table death counts for ages from 0 to 109 in single years of age, with the last age group being age 110+. A distinctive advantage of life-table death counts lies in the availability of data observed across all ages.
@@
Figure 1: Functional time series graphical displays of age-specific life-table death counts from 1975 to 2023 in a single-year group. The oldest years are shown in red, with the most recent years in violet. Curves are ordered chronologically according to the colors of the rainbow.
Figure 1 shows rainbow plots of the female and male age-specific life-table death counts in Japan from 1975 to 2023. The time ordering of the curves follows the color order of a rainbow, where curves from the distant past are shown in red, and the more recent curves are shown in violet. The figures show typical mortality curves for a developed country, with a declining trend in infant mortality. A typical negatively skewed distribution of life-table death counts is apparent, with the peaks shifting to higher ages for both females and males. This shift is a major source of longevity risk, which is a major issue for insurers and pension funds, especially in the sale and risk management of annuity products (see Denuit et al., 2007, for a discussion). In addition, the spread of the distribution indicates the variability in lifespan. A decrease in variability over time can be directly observed and quantified using the Gini coefficient (Wilmoth & Horiuchi, 1999; Debón et al., 2017).
In Figure 2, we present a map of 47 Japanese prefectures from the northern region, Hokkaido, to the southern region, Okinawa. Due to geography and socio-economic status, the age distributions of death counts tend to be heterogeneous. Differences between prefectures or years within the same prefecture can be visualized using image plots in Section 2.1 and quantified using the functional cross-correlation function in Section 2.2.
Figure 2: A geographical map of 47 prefectures in Japan.
@@ 2.1 Image plots
Another visual perspective of the data is shown in the image plot of Figure 3. To measure the distance between two probability densities, such as the subnational and national age distributions of death counts, we consider the symmetric KLD (see Edwards & Tuljapurkar, 2005; d’Albis et al., 2014, for examples of using the KLD to compare country-specific lifespan distributions). For a given prefecture and gender, the symmetric KLD can be expressed as
$KLD_{s}^{g}=\$$D_{KL}\left[d_{t,s}^{g}(u)|d_{t,\text{national}}^{g}(u)\right]+D_{\text{KL}}\left[d_{t,\text{national}}^{g}(u)|d_{t,s}^{g}(u)\right]$
$=\$$\frac{1}{111\times n}\sum^{n}_{t=1}\sum_{i=1}^{111}d_{t,s}^{g}(u_{i})\times[\ln d_{t,s}^{g}(u_{i})-\ln d_{t,\text{national}}^{g}(u_{i})]+$
$\frac{1}{111\times n}\sum^{n}_{t=1}\sum_{i=1}^{111}d_{t,\text{national}}^{g}(u_{i})\times[\ln d_{t,\text{national}}^{g}(u_{i})-\ln d_{t,s}^{g}(u_{i})],$
where $n$ denotes the number of years in the data set. Here, we graph the symmetric KLD between life-table death counts for each prefecture to life-table death counts for the whole country, thus allowing relative mortality comparisons to be made.
A divergent color palette is used, with red representing larger KLD values (i.e., greater difference between subnational and national data) and yellow denoting smaller KLD values. The prefectures are ordered geographically from north to south, so the most northerly prefecture (Hokkaido) is at the top of the panels, and the most southerly prefecture (Okinawa) is at the bottom.
Figure 3: Image plots showing the symmetric KLD between life-table death counts of a prefecture and life-table death counts of the whole country. The top panel shows mortality averaged over ages, while the bottom panel shows mortality averaged over years. Prefectures are numbered geographically from north to south.
The top row of the panels shows mortality for each prefecture and year, averaged over all ages. In 2011, in Iwate, Miyagi, and Fukushima, there was a large increase in mortality compared to other prefectures. These are northern coastal regions, and the elevated relative mortality is due to the tsunami of 11 March 2011 (see, e.g., Nakahara & Ichikawa, 2013) caused by a magnitude–9.0 earthquake.
@@
The bottom of the panels shows the mortality for each prefecture and age, averaged over all years. Several striking features are apparent. There are differences between prefectures in the care of children and adolescents; this may be due to socio-economic differences and the accessibility of health services (Tanaka et al., 2023; Okamoto et al., 2025).
@@ 2.2 Functional cross-correlation plot
Let $d_{t,s}^{g}(u)$ be the age distribution of deaths for age $u$, prefecture $s$, gender $g$ at year $t$, respectively. There are 111 ages, $S=47$ prefectures, and $49$ years. Following Kokoszka et al. (2017) and Mestre et al. (2021), the functional autocorrelation coefficient at lag $h$ can be defined as
$\|C_{h}\|$ $=\int\int C_{h}^{2}(u,u)dudu$
$\rho_{h}$ $=\frac{\|C_{h}\|}{\int C_{0}(u,u)du},$ (1)
@@
$\left\|C_{h}[d_{s}^{g}(u),d_{\text{national}}^{g}(v)]\right\|=\int\int C_{h}^{2}[d_{s}^{g}(u),d_{\text{national}}^{g}(v)]dudv,$
@@
$\rho_{h}[d_{s}^{g}(u),d^{g}_{\text{national}}(v)]=\frac{\left\|C_{h}[d_{s}^{g}(u),d_{\text{national}}^{g}(v)]\right\|}{\sqrt{\int C_{0}\big[d^{g}_{s}(u),d^{g}_{s}(u)\big]du}\sqrt{\int C_{0}\big[d^{g}_{\text{national}}(v),d^{g}_{\text{national}}(v)\big]dv}}.$ (2)
@@
In Figure 4, we display the functional cross-correlation functions showing the dependence of life-table death counts between each prefecture and the whole country. Each subnational data set shows a stronger cross-correlation with the national data at lags  1 and  2. The cross-correlation reduces as lag increases.
@@ 3 Cumulative distribution function transformation
To model life-table death counts, there exists a range of transformations, including centered log-ratio transformation (Aitchison, 1986), $\alpha$-transformation (Shang & Haberman, 2025c), and CDF transformation (Shang & Haberman, 2025b). Among the three, the CDF transformation can handle not only zero counts but also the monotonicity constraint. In the CDF transformation, we first convert an age distribution of death for each year to a probability by dividing its radix $10^{5}$. Via the cumulative sum, we transform a probability density function into a CDF,
@@
where $D_{t,s}^{g}(u_{111})=1$. In doing so, it enjoys the additional benefit of monotonicity (see also Mayhew & Smith, 2013).
@@
where $\Delta$ represents the first-order differencing, where $\Delta_{i=1}^{1}D_{t,s}^{g}(u_{1})=D_{t,s}^{g}(u_{1})$. The constant $10^{5}$ is the radix of the life-table death count. Since the data are observed discretely, we use cumulative sums and first-order differences to approximate integration and differentiation, respectively.
@@ 4 Functional time-series forecasting models
We obtain a set of unconstrained functional time series $(\mathcal{X}_{1,s}^{g}(u),\dots,\mathcal{X}_{n,s}^{g}(u)]$ via the CDF transformation. Generally, we represent it as $\{\mathcal{X}_{t,s}^{g}(u)\}$, which is an arbitrary functional time series defined in a common probability space. It is assumed that the observations $\mathcal{X}_{t,s}^{g}(u)$ are elements of the Hilbert space $\mathcal{H}=\mathcal{L}^{2}(\mathcal{I})$ equipped with the inner product $\langle w,v\rangle=\int_{\mathcal{I}}w(u)v(u)du$, where $u$ represents a continuum and $\mathcal{I}\subset\mathds{R}$ denotes a function support range and $\mathds{R}$ is the real line. Each function is a square-integrable function that satisfies $\|\mathcal{X}_{t,s}^{g}\|^{2}=\int_{\mathcal{I}}(\mathcal{X}_{t,s}^{g}(u))^{2}du<\infty$ and the associated norms.